OpenStudy (anonymous):

Which number line correctly shows 1.2, , and ? A. B. C. D.

jabez177 (jabez177):

http://prntscr.com/9mj9oz <- Use this to take a snapshot of it. :)

jabez177 (jabez177):

Then you can take a screenshot and paste the link here and I will be able to answer it. :)

OpenStudy (anonymous):

http://prntscr.com/9mjbvl

OpenStudy (anonymous):

can you pls explain how to solve

OpenStudy (anonymous):

are you still here

OpenStudy (anonymous):

@jabez177

jabez177 (jabez177):

Yes, I'm back!

jabez177 (jabez177):

So all of those number lines are in marks of tens. So 1.2 is basically \[1\frac{ 2 }{ 10 }\]

jabez177 (jabez177):

\[1\frac{ 3 }{ 5 }=1\frac{ 6 }{ 10 }\]

jabez177 (jabez177):